# Retraining with Predicted Hard Labels Provably Increases Model Accuracy

### v1

@@ Appendix D Proof of Theorem 3
$\geq\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$
$\geq\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}(\widehat{I}\neq i|I=i)$ (92)
$=\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\mathbb{P}(\widehat{I}\neq I)\,,$ (93)
@@
with $\Phi$ denoting the CDF of a standard normal variable. Equation 92 above follows from the lemma below.
Lemma 5.
@@
$\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)\geq\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\;\mathbb{P}(\widehat{I}\neq i|I=i)\,.$
Combining (91) and (93), we obtain that
$\delta_{0}:=\frac{\delta}{\Phi\Big{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Big{)}}\geq\mathbb{P}(\widehat{I}\neq I)\,.$ (94)
Next recall the set $\mathcal{T}:=\{({\mathbf{x}}_{j},\widehat{\mathrm{y}}_{j})\}_{j\in[n]}$, and let $\mathbf{X}=[{\mathbf{x}}_{1}^{T},\dotsc,{\mathbf{x}}_{n}^{T}]^{T}\in\mathbb{R}^{n\times d}$ and $\widehat{\mathbf{y}}=[\widehat{\mathrm{y}}_{1},\dotsc,\widehat{\mathrm{y}}_{n}]^{T}$. By an application of Fano’s inequality, with conditioning on $\mathbf{X}$ (see e.g. (Scarlett and Cevher,, 2019, Section 2.3)) we have
$\mathcal{I}(I;\widehat{I}|\mathbf{X})\geq\left(1-\delta_{0}\right)\log(M(\rho,S^{d-1}))-\log 2\,,$ (95)
where $\mathcal{I}(I;\widehat{I}|\mathbf{X})$ represents the conditional mutual information between $I$ and $\widehat{I}$. Using the fact that $I\to\widehat{\mathbf{y}}\to\widehat{I}$ forms a Markov chain conditioned on $\mathbf{X}$, and by an application of the data processing inequality we have:
$\mathcal{I}(I;\widehat{I}|\mathbf{X})\leq\mathcal{I}(I;\widehat{\mathbf{y}}|\mathbf{X})$ (96)
We will now upper bound $\mathcal{I}(I;\widehat{\mathbf{y}}|\mathbf{X})$. Let $\mathrm{w}_{j}:=\frac{1}{2}\big{(}\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)+1\big{)}$ be the 0-1 version of the actual label of $\mathbf{x}_{j}$, viz., $\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)$. As per our setting, we have $\widehat{\mathrm{y}}_{j}=2\big{(}\mathrm{w}_{j}\oplus z_{j}\big{)}-1$ where $z_{j}\sim\text{Bernoulli}(p)$ and $\oplus$ denotes modulo-2 addition. Since the noise variables $z_{j}$ are independent and $\widehat{\mathrm{y}}_{j}$ depends on $(I,\mathbf{X})$ only through $\mathrm{w}_{j}=\frac{1}{2}\big{(}\mathrm{sign}(\langle\mathbf{x}_{j},\bm{\theta}_{I}\rangle)+1\big{)}$, by using the tensorization property of the mutual information (see e.g. (Scarlett and Cevher,, 2019, Lemma 2, part (iii))), we have
$\mathcal{I}(I;\mathbf{y}|\mathbf{X})\leq\sum_{j=1}^{n}\mathcal{I}\big{(}\mathrm{w}_{j};\widehat{\mathrm{y}}_{j}\big{)}\leq n(\log 2-H_{2}(p))\,,$ (97)
where the second inequality follows since $\widehat{\mathrm{y}}_{j}$ is generated by passing $\mathrm{w}_{j}$ through a binary symmetric channel, which has capacity $\log 2-H_{2}(p)$ with $H_{2}(p):=-p\log p-(1-p)\log(1-p)$ denoting the binary entropy function.
We next use the lemma below to further upper bound the right-hand side of (97).
Lemma 6.
@@
Using Lemma 6 with $k=2$ we obtain $\log 2-H_{2}(p)\leq 4(p-1/2)^{2}=(1-2p)^{2}$, which along with (97), (96) and (95) gives
$\frac{n(1-2p)^{2}+\log 2}{1-\delta_{0}}\geq\log(M(\rho,S^{d-1}))\,.$ (98)
In our next lemma, we lower bound $M(\rho,S^{d-1})$.
Lemma 7.
@@
Recall the definition of $\rho$-packing number of $S^{d-1}$ given by (90). We have the following bound:
$M(\rho,S^{d-1})\geq\exp\Big{(}\frac{d\rho^{2}}{2}\Big{)}\,.$
@@
Using Lemma 7 along with (98), we obtain the following lower bound on the sample complexity:
$n\geq\frac{\frac{\rho^{2}}{2}(1-\delta_{0})d-\log 2}{(1-2p)^{2}},\text{ with }\delta_{0}=\frac{\delta}{\Phi\Big{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Big{)}}.$
Let us pick $\rho=\frac{(4/5)^{2}-\mu^{2}}{(4/5)^{2}+\mu^{2}}$; it can be shown that $\rho>\frac{2}{5}$, because $\mu\leq\frac{1}{2}$ as per our setting. In that case, using the fact that $\Phi(-\frac{4}{5})>\frac{1}{5}$, we have that $\delta_{0}<5\delta$. Using all of this gives us the desired result for $\delta<\frac{1}{5}$.
We now prove Lemmas 5, 6 and 7.
@@
Proof of Lemma 5. Let us consider the event $\widehat{I}\neq i$, given that $I=i$. We note that if $\widehat{I}\neq I$, then $\langle\widehat{\bm{\theta}},\bm{\theta}_{I}\rangle\leq\sqrt{\frac{1+\rho}{2}}$ or equivalently, the angle between $\widehat{\bm{\theta}}$ and $\bm{\theta}_{I}$ is $\geq b:=\cos^{-1}\Big{(}\sqrt{\frac{1+\rho}{2}}\Big{)}$. Otherwise, by definition of $\widehat{I}$ we have $\langle\widehat{\bm{\theta}},\bm{\theta}_{\widehat{I}}\rangle\geq\langle\widehat{\bm{\theta}},\bm{\theta}_{I}\rangle>\sqrt{\frac{1+\rho}{2}}$, and therefore the angle between $\bm{\theta}_{\widehat{I}}$ and $\bm{\theta}_{I}$ is $<2b$. Noting that $\cos(2b)=2\cos^{2}(b)-1=\rho$, we would then have $\langle\bm{\theta}_{\widehat{I}},\bm{\theta}_{I}\rangle>\rho$, which is a contradiction since $\Theta$ forms a $\rho$-packing.
Next, under our data model and conditioned on the value of $I$, $\mathbb{E}\big{[}|\langle\mathbf{x},\bm{\theta}_{I}\rangle|\big{]}=\mu$ and on the space orthogonal to $\bm{\theta}_{I}$, $\mathbf{x}$ follows $\mathcal{N}(0,\sigma^{2}\mathbf{I}_{d-1})$ distribution. By rotating the basis and without loss of generality in lower bounding $\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$, we can assume $\bm{\theta}_{I}=\bm{\theta}_{i}=\mathbf{e}_{1}$ and $\widehat{\bm{\theta}}$ is an arbitrary vector in $S^{d-1}$ such that $\theta:=\langle\widehat{\bm{\theta}},\bm{e}_{1}\rangle\leq\sqrt{\frac{1+\rho}{2}}$. We proceed by writing
@@
$\mathbb{P}\left(\mathrm{sign}(\langle\mathbf{x},\bm{\theta}_{I}\rangle)\neq\mathrm{sign}(\langle\mathbf{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)=\mathbb{P}\left(\mathrm{sign}({\mathrm{x}}^{(1)})\Big{(}{\mathrm{x}}^{(1)}\theta+\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\Big{)}\leq 0\right)\,,$
where for a vector ${\mathbf{v}}\in\mathbb{R}^{d}$, we use the notation ${\mathbf{v}}^{(-1)}\in\mathbb{R}^{d-1}$ to refer to the vector obtained by dropping its first coordinate. Note that $\mathbf{x}$ is a test data point, independent of the training data $\mathcal{T}$ and so it is independent of $\bm{\widehat{\theta}}$. Therefore, we have $\mathrm{sign}({\mathrm{x}}^{(1)})\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\sim\mathcal{N}(0,\sigma^{2}\|\widehat{\bm{\theta}}^{(-1)}\|^{2})$. By invoking the condition $\|\widehat{\bm{\theta}}\|=1$, this can be written as $\mathrm{sign}({\mathrm{x}}^{(1)})\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle=\sqrt{1-\theta^{2}}\,Z$, where $Z$ is a standard normal variable. Hence, we have:
@@
$\mathbb{P}\left(\mathrm{sign}({\mathrm{x}}^{(1)})\Big{(}{\mathrm{x}}^{(1)}\theta+\langle\mathbf{x}^{(-1)},\widehat{\bm{\theta}}^{(-1)}\rangle\Big{)}\leq 0\right)$
@@
$=\mathbb{P}\left(|{\mathrm{x}}^{(1)}|\theta+\sqrt{1-\theta^{2}}\,Z\leq 0\right)$
@@
$=\mathbb{P}\left(Z\leq-\frac{|{\mathrm{x}}^{(1)}|\theta}{\sqrt{1-\theta^{2}}}\right)$
$=1-\mathbb{E}\Bigg{[}\Phi\Big{(}\frac{\theta}{\sqrt{1-\theta^{2}}}|{\mathrm{x}}^{(1)}|\Big{)}\Bigg{]}$
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}1-\mathbb{E}\Bigg{[}\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}|{\mathrm{x}}^{(1)}|\Bigg{)}\Bigg{]}$
@@
$\stackrel{{\scriptstyle(b)}}{{\geq}}1-\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}\mathbb{E}\big{[}|{\mathrm{x}}^{(1)}|\big{]}\Bigg{)}$
$=1-\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}$
$=\Phi\Bigg{(}-\sqrt{\frac{1+\rho}{1-\rho}}\mu\Bigg{)}\,,$
where $(a)$ follows from the fact that $\theta\leq\sqrt{\frac{1+\rho}{2}}$ and $(b)$ holds due to Jensen’s inequality and concavity of $\Phi(.)$ on the positive values. Combining the above bound with the event $\widehat{I}\neq i$ given that $I=i$ gives us the desired result.
@@
Proof of Lemma 6. Define $q_{i}=p_{i}-1/k$. Note that $q_{i}$ can be negative, and we have $\sum_{i=1}^{k}q_{i}=0$. We write
@@
Note that in eq. 99 we used the fact that $1+kq_{i}\geq 0$ and $\log x\leq x-1$ for all $x\geq 0$.

### v2

@@ B.2 Lower bound $\widetilde{\alpha}_{0}(\mathbf{x})$
$=1-\frac{1}{2}\operatorname{\mathbb{E}}\left[\exp\left(-\frac{n(1-2p)^{2}\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{8\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$ (C.6)
Under our data model, $\bm{x}=y(1+u){\bm{{\mu}}}+\bm{\Sigma}^{1/2}{\bm{z}}$. Consider the probabilistic event $\mathcal{E}_{2}:=\{{\bm{z}}:\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2d\}$. For every $\eta>0$, we have
$\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]$ $=\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)(\mathbb{1}_{\mathcal{E}_{2}}+\mathbb{1}_{\mathcal{E}_{2}^{c}})\right]$
$\leq\exp\left(-\frac{\eta\gamma^{4}}{2\lambda_{\max}d}\right)+\mathbb{P}(\mathcal{E}_{2}^{c})\,,$
where the inequality holds because $\langle\bm{x},{\bm{{\mu}}}\rangle=y(1+u)\gamma^{2}$ due to which $|\langle\bm{x},{\bm{{\mu}}}\rangle|\geq\gamma^{2}$ (given that $u$ is a non-negative random variable) and because,
$\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}=\left\|\bm{\Sigma}{\bm{z}}\right\|_{\ell_{2}}^{2}\leq\lambda_{\max}^{2}\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\leq 2\lambda_{\max}^{2}d$
when event $\mathcal{E}_{2}$ happens. Further, $\left\|{\bm{z}}\right\|_{\ell_{2}}^{2}\sim\chi^{2}_{d}$ and so using tail bounds for $\chi^{2}_{d}$, see e.g., (Ghosh,, 2021, Theorem 3), we have $\mathbb{P}(\mathcal{E}_{2}^{c})\leq e^{-d/8}$. Putting things together, we obtain
$\operatorname{\mathbb{E}}\left[\exp\left(-\frac{\eta\langle\bm{x},{\bm{{\mu}}}\rangle^{2}}{\left\|\bm{\Sigma}^{1/2}\bm{x}\right\|_{\ell_{2}}^{2}}\right)\right]\leq\exp\left(-\frac{\eta\gamma^{4}}{2\lambda_{\max}^{2}d}\right)+e^{-d/8}\,.$
Setting $\eta=n(1-2p)^{2}/8$ and using this in (C.6), we get
${\sf acc}({\widehat{\bm{\theta}}}_{0})\geq 1-\frac{1}{2}\exp\left(-\frac{n(1-2p)^{2}\gamma^{4}}{16\lambda_{\max}^{2}d}\right)-e^{-d/8}-2\exp\left(-\frac{n}{32}(1-2p)^{2}\right)\,.$
Appendix D Proof of Theorem 4.6
Proof Since we are interested with dependence of sample size with respect to $d$ and $p$, we assume $\gamma:=\left\|{\bm{{\mu}}}\right\|_{\ell_{2}}=1$ and $\bm{\Sigma}=\bm{I}-{\bm{{\mu}}}{\bm{{\mu}}}^{\sf T}$ the projection onto the orthogonal space of ${\bm{{\mu}}}$. Also note that without loss of generality, we can assume that our estimator $\widehat{\bm{\theta}}$ has unit norm, because its norm does not affect the sign of $\langle\bm{x},{\widehat{\bm{\theta}}}\rangle$. This way, ${\bm{{\mu}}}$ and ${\widehat{\bm{\theta}}}$ both belong to $\mathbb{S}^{d-1}$.
We first follow a standard argument to “reduce” the classification problem to a multi-way hypothesis testing problem. Let $\rho\in(0,1)$ be an arbitrary but fixed value which we can choose. We define a $\rho$-packing of $\mathbb{S}^{d-1}$ as a set =$\mathcal{M}=\{{\bm{{\mu}}}_{1},\dotsc,{\bm{{\mu}}}_{M}\}\subset\mathbb{S}^{d-1}$ such that $\langle{\bm{{\mu}}}_{l},{\bm{{\mu}}}_{k}\rangle\leq\rho$ for $l\neq k$. Also define the $\rho$-packing number of $\mathbb{S}^{d-1}$ as
$M(\rho,\mathbb{S}^{d-1}):=\sup\{M\in\mathbb{N}:\text{there exists a $\rho$-packing $\mathcal{M}$ of $\mathbb{S}^{d-1}$ with size $M$}\}\,.$ (D.1)
For convenience, we define the misclassification error of a classifier $\widehat{\bm{\theta}}\in\mathbb{S}^{d-1}$ as $\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})=1-\textup{acc}(\widehat{\bm{\theta}};{\bm{{\mu}}})$. As per our condition,
$\delta\geq\sup_{{\bm{{\mu}}}\in\mathbb{S}^{d-1}}\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})\geq\sup_{{\bm{{\mu}}}\in\mathcal{M}}\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})\,.$ (D.2)
In order to further lower bound the right hand side, we let $I$ be a random variable uniformly distributed on the hypothesis set $\{1,2,\dotsc,M\}$ and consider the case of ${\bm{{\mu}}}={\bm{{\mu}}}_{I}$. We also define $\widehat{I}$ as the index of the element in $\mathcal{M}$ with maximum inner product with $\widehat{\bm{\theta}}$ (it does not matter how we break ties). Under our data model we have $\langle\bm{x},{\bm{{\mu}}}_{I}\rangle=y(1+u)\left\|{\bm{{\mu}}}_{I}\right\|_{\ell_{2}}^{2}=y(1+u)$ and since $u$ is a non-negative random variable, $y={\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)$. We then have
$\sup_{{\bm{{\mu}}}\in\mathcal{M}}\textup{err}(\widehat{\bm{\theta}};{\bm{{\mu}}})$ $\geq\max_{i\in[M]}\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$
$\geq\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$
$\geq\Phi\Bigg{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Bigg{)}\frac{1}{M}\sum_{i=1}^{M}\mathbb{P}(\widehat{I}\neq i|I=i)$ (D.3)
$=\Phi\Bigg{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Bigg{)}\mathbb{P}(\widehat{I}\neq I)\,,$ (D.4)
@@
with $\Phi$ denoting the CDF of a standard normal variable. Equation D.3 above follows from the lemma below.
Lemma D.1
@@
$\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)\geq\Phi\Bigg{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Bigg{)}\;\mathbb{P}(\widehat{I}\neq i|I=i)\,.$
Combining (D.2) and (D.4), we obtain that
$\delta_{0}:=\frac{\delta}{\Phi\Big{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Big{)}}\geq\mathbb{P}(\widehat{I}\neq I)\,.$ (D.5)
Next recall the set $\mathcal{T}:=\{(\bm{x}_{j},\widehat{y}_{j})\}_{j\in[n]}$, and let ${\bm{X}}=[\bm{x}_{1}^{T},\dotsc,\bm{x}_{n}^{T}]^{T}\in\mathbb{R}^{n\times d}$ and $\widehat{\bm{y}}=[\widehat{y}_{1},\dotsc,\widehat{y}_{n}]^{T}$. By an application of Fano’s inequality, with conditioning on $\mathbf{X}$ (see e.g. (Scarlett and Cevher,, 2019, Section 2.3)) we have
$\mathcal{I}(I;\widehat{I}|{\bm{X}})\geq\left(1-\delta_{0}\right)\log(M(\rho,\mathbb{S}^{d-1}))-\log 2\,,$ (D.6)
where $\mathcal{I}(I;\widehat{I}|{\bm{X}})$ represents the conditional mutual information between $I$ and $\widehat{I}$. Using the fact that $I\to\widehat{\bm{y}}\to\widehat{I}$ forms a Markov chain conditioned on $\mathbf{X}$, and by an application of the data processing inequality we have:
$\mathcal{I}(I;\widehat{I}|{\bm{X}})\leq\mathcal{I}(I;\widehat{\bm{y}}|{\bm{X}})$ (D.7)
We will now upper bound $\mathcal{I}(I;\widehat{\bm{y}}|{\bm{X}})$. Let $w_{j}:=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$ be the 0-1 version of the actual label of $\bm{x}_{j}$, viz., ${\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)$. As per our setting, we have $\widehat{y}_{j}=2\big{(}w_{j}\oplus z_{j}\big{)}-1$ where $z_{j}\sim\text{Bernoulli}(p)$ and $\oplus$ denotes modulo-2 addition. Since the noise variables $z_{j}$ are independent and $\widehat{y}_{j}$ depends on $(I,{\bm{X}})$ only through $w_{j}=\frac{1}{2}\big{(}{\rm sign}(\langle\bm{x}_{j},{\bm{{\mu}}}_{I}\rangle)+1\big{)}$, by using the tensorization property of the mutual information (see e.g. (Scarlett and Cevher,, 2019, Lemma 2, part (iii))), we have
$\mathcal{I}(I;\bm{y}|{\bm{X}})\leq\sum_{j=1}^{n}\mathcal{I}\big{(}w_{j};\widehat{y}_{j}\big{)}\leq n(\log 2-H_{2}(p))\,,$ (D.8)
where the second inequality follows since $\widehat{y}_{j}$ is generated by passing $w_{j}$ through a binary symmetric channel, which has capacity $\log 2-H_{2}(p)$ with $H_{2}(p):=-p\log p-(1-p)\log(1-p)$ denoting the binary entropy function.
We next use the lemma below to further upper bound the right-hand side of (D.8).
Lemma D.2
@@
Using Lemma D.2 with $k=2$ we obtain $\log 2-H_{2}(p)\leq 4(p-1/2)^{2}=(1-2p)^{2}$, which along with (D.8), (D.7) and (D.6) gives
$\frac{n(1-2p)^{2}+\log 2}{1-\delta_{0}}\geq\log(M(\rho,\mathbb{S}^{d-1}))\,.$ (D.9)
In our next lemma, we lower bound $M(\rho,\mathbb{S}^{d-1})$.
Lemma D.3
@@
Recall the definition of $\rho$-packing number of $\mathbb{S}^{d-1}$ given by (D.1). We have the following bound:
$M(\rho,\mathbb{S}^{d-1})\geq\exp\Big{(}\frac{d\rho^{2}}{2}\Big{)}\,.$
@@
Using Lemma D.3 along with (D.9), we obtain the following lower bound on the sample complexity:
$n\geq\frac{\frac{\rho^{2}}{2}(1-\delta_{0})d-\log 2}{(1-2p)^{2}},\text{ with }\delta_{0}=\frac{\delta}{\Phi\Big{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Big{)}}.$
Note that $\rho\in(0,1)$ can be set arbitrarily. Since our claim is on the order of sample size, the specific value of $\rho$ does not matter.
We now prove Lemmas D.1, D.2 and D.3.
@@
Proof of Lemma D.1. Let us consider the event $\widehat{I}\neq i$, given that $I=i$. We note that if $\widehat{I}\neq I$, then $\langle\widehat{\bm{\theta}},{\bm{{\mu}}}_{I}\rangle\leq\sqrt{\frac{1+\rho}{2}}$ or equivalently, the angle between $\widehat{\bm{\theta}}$ and ${\bm{{\mu}}}_{I}$ is $\geq b:=\cos^{-1}\Big{(}\sqrt{\frac{1+\rho}{2}}\Big{)}$. Otherwise, by definition of $\widehat{I}$ we have $\langle\widehat{\bm{\theta}},{\bm{{\mu}}}_{\widehat{I}}\rangle\geq\langle\widehat{\bm{\theta}},{\bm{{\mu}}}_{I}\rangle>\sqrt{\frac{1+\rho}{2}}$, and therefore the angle between ${\bm{{\mu}}}_{\widehat{I}}$ and ${\bm{{\mu}}}_{I}$ is $<2b$. Noting that $\cos(2b)=2\cos^{2}(b)-1=\rho$, we would then have $\langle{\bm{{\mu}}}_{\widehat{I}},{\bm{{\mu}}}_{I}\rangle>\rho$, which is a contradiction since $\mathcal{M}$ forms a $\rho$-packing. We proceed by writing
@@
$\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i\right)$ $\geq\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i,\widehat{I}\neq i\right)\mathbb{P}(\widehat{I}\neq i|I=i)\,.$
As discussed above, on the event that $I=i,\widehat{I}\neq i$, we have $\theta:=\langle{\widehat{\bm{\theta}}},{\bm{{\mu}}}_{i}\rangle\leq\sqrt{\frac{1+\rho}{2}}$. Consider the decomposition $\bm{x}=\langle\bm{x},{\bm{{\mu}}}_{i}\rangle{\bm{{\mu}}}_{i}+\mathcal{P}^{\perp}_{{\bm{{\mu}}}_{i}}\bm{x}$. We then have
@@
$\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i,\widehat{I}\neq i\right)=\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{i}\rangle)\left(\langle\bm{x},{\bm{{\mu}}}_{i}\rangle\theta+\mathcal{P}^{\perp}_{\bm{\mu}_{i}}\bm{x},\mathcal{P}^{\perp}_{\bm{\mu}_{i}}{\widehat{\bm{\theta}}}\rangle\right)\leq 0\Big{|}I=i,\widehat{I}\neq i\right)\,.$
Note that $\bm{x}$ is a test data point, independent of the training data $\mathcal{T}$ and so it is independent of ${\widehat{\bm{\theta}}}$. In addition, under our data model, $\langle\bm{x},{\bm{{\mu}}}_{i}\rangle$ is independent of $\mathcal{P}^{\perp}_{\bm{\mu}_{i}}\bm{x}$. Hence, ${\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{i}\rangle)\langle\mathcal{P}^{\perp}_{\bm{\mu}_{i}}\bm{x},\mathcal{P}^{\perp}_{{\bm{{\mu}}}_{i}}{\widehat{\bm{\theta}}}\rangle\sim{\mathcal{N}}(0,\left\|\mathcal{P}^{\perp}_{\bm{\mu}_{i}}{\widehat{\bm{\theta}}}\right\|_{\ell_{2}}^{2})$. Given that $\left\|{\widehat{\bm{\theta}}}\right\|_{\ell_{2}}=1$ we also have $\left\|\mathcal{P}^{\perp}_{\bm{\mu}_{i}}{\widehat{\bm{\theta}}}\right\|_{\ell_{2}}^{2}=1-\langle{\widehat{\bm{\theta}}},{\bm{{\mu}}}_{i}\rangle^{2}=1-\theta^{2}$. In short, we can write
${\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{i}\rangle)\langle\mathcal{P}^{\perp}_{{\bm{{\mu}}}_{i}}\bm{x},\mathcal{P}^{\perp}_{{\bm{{\mu}}}_{i}}{\widehat{\bm{\theta}}}\rangle=\sqrt{1-\theta^{2}}Z,\quad Z\sim{\mathcal{N}}(0,1)\,.$
Using this characterization, we proceed by writing
@@
$\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{I}\rangle)\neq{\rm sign}(\langle\bm{x},\widehat{\bm{\theta}}\rangle)\Big{|}I=i,\widehat{I}\neq i\right)$ $=\mathbb{P}\left({\rm sign}(\langle\bm{x},{\bm{{\mu}}}_{i}\rangle)\langle\bm{x},{\bm{{\mu}}}_{i}\rangle\theta+\sqrt{1-\theta^{2}}Z\leq 0\right)$
@@
$=\mathbb{P}\left(Z\leq-\frac{|\langle\bm{x},{\bm{{\mu}}}_{i}\rangle|\theta}{\sqrt{1-\theta^{2}}}\right)$
$=1-\mathbb{E}\Bigg{[}\Phi\Big{(}\frac{\theta}{\sqrt{1-\theta^{2}}}|\langle\bm{x},{\bm{{\mu}}}_{i}\rangle|\Big{)}\Bigg{]}$
@@
$\stackrel{{\scriptstyle(a)}}{{\geq}}1-\mathbb{E}\Bigg{[}\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}|\langle\bm{x},{\bm{{\mu}}}_{i}\rangle|\Bigg{)}\Bigg{]}$
@@
$\stackrel{{\scriptstyle(b)}}{{\geq}}1-\Phi\Bigg{(}\sqrt{\frac{1+\rho}{1-\rho}}\mathbb{E}\big{[}|\langle\bm{x},{\bm{{\mu}}}_{i}\rangle|\big{]}\Bigg{)}$
$\stackrel{{\scriptstyle(c)}}{{\geq}}1-\Phi\Bigg{(}2\sqrt{\frac{1+\rho}{1-\rho}}\Bigg{)}=\Phi\Bigg{(}-2\sqrt{\frac{1+\rho}{1-\rho}}\Bigg{)}\,,$
where $(a)$ follows from the fact that $\theta\leq\sqrt{\frac{1+\rho}{2}}$; $(b)$ holds due to Jensen’s inequality and concavity of $\Phi(\cdot)$ on the positive values, and $(c)$ holds because under our data model $\operatorname{\mathbb{E}}[|\langle\bm{x},{\bm{{\mu}}}_{i}\rangle|]=\operatorname{\mathbb{E}}[y(1+u)\left\|{\bm{{\mu}}}_{i}\right\|_{\ell_{2}}^{2}]=\operatorname{\mathbb{E}}[1+u]\leq 2$ since $\|u\|_{\psi_{2}}=1$. This completes the proof of claim.
@@
Proof of Lemma D.2. Define $q_{i}=p_{i}-1/k$. Note that $q_{i}$ can be negative, and we have $\sum_{i=1}^{k}q_{i}=0$. We write
@@
Note that in eq. D.10 we used the fact that $1+kq_{i}\geq 0$ and $\log x\leq x-1$ for all $x\geq 0$.